OpenStudy (anonymous):

2i/3-4i write as a complex number in standard form

OpenStudy (anonymous):

Standard form = a+ib

OpenStudy (anonymous):

\[\frac{2i}{3-4i}\times \frac{3+4i}{3+4i}\] does it and this works because \[(a+bi)(a-bi)=a^2+b^2\] a real number

OpenStudy (anonymous):

Rationalise the denominator as we don't want i in the denominator